OpenStudy (anonymous):

Solve the equation. x2 + 8x + 16 = 49

OpenStudy (anonymous):

10x+16

OpenStudy (atlas):

x2 + 8x + 16 - 49 =0; x2 + 8x - 33 =0; x2 +11x-3x -33 =0; x(x+11) -3(x+11) =0; (x-3)(x+11) =0; x =3; x=-11

OpenStudy (atlas):

the trick is to know how to break the middle term to factorize the quadratic equation

OpenStudy (atlas):

yea u are right

hero (hero):

\(x^2 + 8x + 16) = 49\) \((x + 4)^2 = 7^2\) Square root both sides: \(x + 4 = \pm 7\) \(x = \pm 7 - 4\)